OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b?

OpenStudy (kamibug):

\(\huge\bf\color{green}{Welcome~to~Openstudy!~:)}\) a - b = 12 5a - 2b = 18 a = 12 + b 5(12+b) - 2b = 18 60 + 5b - 2b = 18 60 + 3b = 18 3b = -42 b = -14 <--- a -(-14) = 12 a + 14 = 12 a = -2 <---